You are eyeing an investment in Treasury Notes for a full lot (i.e. PAR=$100,000), and find a note with a maturity of 20 years. This particular Treasury has a YTM of 4.16%, and a stated rate of interest of 0.89% with a maturity of 20 years. The market price of this treasury is closest to

a. $50,219.77

b. $55,894.94

c. $56,182.67

d. $67,899.07
Answer #1

Coupon = 0.0089 * 100,000 = 890

Price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Price = 890 * [1 - 1 / (1 + 0.0416)20] / 0.0416 + 100,000 / (1 + 0.0416)20

Price = 890 * [1 - 0.442569] / 0.0416 + 44,256.85774

Price = 890 * 13.399794 + 44,256.85774

Price = $56,182.67